OpenStudy (anonymous):

12xy(x+y), 12x^2y+12xy^2 multiply

OpenStudy (inkyvoyd):

This ain't healthy man. We are here to help you learn, not do your homework.

OpenStudy (anonymous):

I am not asking you to do my homework just how to do this problem you do not have to be nasty, you will be reported.

hartnn (hartnn):

12x^2y+12xy^2 = 12xy(x+y) so,its like you are squaring that expression.

OpenStudy (inkyvoyd):

sorry @cdelomas - I saw the question you closed and I assumd there were more - I checked and they're weren't. My apologies for misunderstanding.

OpenStudy (anonymous):

thanks how would Id o this problem

OpenStudy (inkyvoyd):

Like hartnn said, follow the distributive property

OpenStudy (anonymous):

Howwould I square it and use thed istributive property can you show me step by step

hartnn (hartnn):

its like \((pq)^2=p^2q^2 \\ so, [12xy(x+y)]^2=12^2x^2y^2(x+y)^2\) got this ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

now 12^2 =144 do u have choices/options ? if not, you can leave the answer as \(144x^2y^2(x+y)^2\)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^